what is angular displacement? pls explain with example..

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Aditya

The difference between the initial and final angular position of a moving body. It is measured in radians. Angular displacement has both magnitude and direction. Conventionally, clockwise movements are described as positive (+) and anticlockwise movements as negative (−). Also, angular displacements of human body segments usually indicate the type of joint movement (e.g. flexion or extension).

Angular displacement

Angular displacement is measured in radians rather than degrees. This is because it provides a very simple relationship between distance traveled around the circle and the distance r from the centre.

\theta=\frac sr

For example if an object rotates 360 degrees around a circle radius r the angular displacement is given by the distance traveled the circumference which is 2πr Divided by the radius in: \theta= \frac{2\pi r}r which easily simplifies to θ = 2π. Therefore 1 revolution is radians.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..